### v5

# UAVDB: Trajectory-Guided Adaptable Bounding Boxes for UAV Detection

Abstract
The widespread deployment of Unmanned Aerial Vehicles (UAVs) in surveillance, security, and airspace management has created an urgent demand for precise, scalable, and efficient UAV detection. However, existing datasets often suffer from limited scale diversity and inaccurate annotations, hindering robust model development. This paper introduces UAVDB, a high-resolution UAV detection dataset constructed using Patch Intensity Convergence (PIC). This novel technique automatically generates high-fidelity bounding box annotations from UAV trajectory data [15], eliminating the need for manual labeling. UAVDB features single-class annotations with a fixed-camera setup and consists of RGB frames capturing UAVs across various scales, from large-scale UAVs to near-single-pixel representations, along with challenging backgrounds that pose difficulties for modern detectors. We first validate the accuracy and efficiency of PIC-generated bounding boxes by comparing Intersection over Union (IoU) performance and runtime against alternative annotation methods, demonstrating that PIC achieves higher annotation accuracy while being more efficient. Subsequently, we benchmark UAVDB using state-of-the-art (SOTA) YOLO-series detectors, establishing UAVDB as a valuable resource for advancing long-range and high-resolution UAV detection. The source code is available at https://github.com/wish44165/UAVDB.
1 Introduction
Precise UAV detection is critical for effective monitoring and threat response. While modern object detection algorithms, such as YOLO-series detectors [11, 24, 23, 10, 22] and transformer-based models [32, 2], have significantly advanced UAV detection, their performance is highly dependent on high-quality annotations. Without accurate, well-annotated datasets, even SOTA models struggle with real-world UAV detection, particularly for tiny UAVs. Existing UAV datasets can be broadly categorized into two types. The first type is ground-target UAV datasets, where UAV-mounted cameras capture objects like vehicles or pedestrians on the ground [25, 26, 7, 31, 18] and the second type is UAV-target datasets, where the UAV itself is the detection target. The latter can be further divided into three categories: 1) RGB frame with fixed-camera setup, where the camera remains stationary as presented in [21, 17], 2) RGB image with moving-camera setup, where the camera equipped on the UAV such as [19, 14], and 3) Infrared image UAV datasets, including single-frame datasets [6, 4, 5], and video-based Anti-UAV datasets [8, 9, 29, 30] which has been featured in four major challenge events.
Figure 1: UAV trajectory captured by Camera 3 in Dataset 4 at 3840$\times$2160 resolution in [15]. The yellow path represents the UAV’s trajectory. On the left, the UAV appears at a short distance with a size of 166$\times$126 pixels, occupying approximately 0.252% of the total image area. On the right, the UAV is shown at a long distance, with a size of 35$\times$36 pixels, covering approximately 0.015% of the entire image. This figure shows the varying visibility of the UAV depending on its distance from the camera.
[…]
0 5334 / 1920$\times$1080 4377 / 1920$\times$1080 33875 / 1920$\times$1080 31075 / 1920$\times$1080 20970 / 1920$\times$1080
1 4941 / 1920$\times$1080 4749 / 1920$\times$1080 19960 / 1920$\times$1080 15409 / 1920$\times$1080 28047 / 1920$\times$1080
2 8016 / 1920$\times$1080 8688 / 1920$\times$1080 17166 / 3840$\times$2160 15678 / 1920$\times$1080 31860 / 2704$\times$2028
3 4080 / 1920$\times$1080 4332 / 1920$\times$1080 14196 / 1440$\times$1080 10933 / 3840$\times$2160 31992 / 1920$\times$1080
4 18900 / 1920$\times$1080 17640 / 1920$\times$1080 21523 / 2288$\times$1080
5 28080 / 1920$\times$1080 32016 / 1920$\times$1080 17550 / 1920$\times$1080
6 11292 / 1440$\times$1080
Table 1: Summary of dataset characteristics in [15]. The table displays the number of frames and resolution for each camera across different datasets. Each cell lists the number of frames followed by the resolution in pixels.
However, existing RGB frames with fixed-camera setup datasets often contain relatively large UAVs or imprecise bounding box annotations, lacking the scale diversity necessary for robust detection models. To address this, we introduce UAVDB, a high-resolution RGB frame featuring multiscale UAVs designed to improve UAV detection in diverse and complex environments. This dataset is particularly relevant for monitoring incoming UAVs from buildings or national borders using a fixed-camera setup. To construct UAVDB, we propose PIC, a technique that automatically generates accurate bounding boxes from trajectory data in [15]. Since their dataset primarily focuses on 3D UAV trajectory reconstruction with unsynchronized consumer cameras and unknown viewpoints, it lacks the precise bounding box annotations required for object detection. Fig. 1 illustrates UAV trajectories alongside human-labeled bounding boxes at different scales, highlighting the need for precise annotations. A detailed dataset structure is provided in Tab. 1. Our contributions are as follows:
  1. $1$
    Introduce UAVDB, a high-resolution RGB frame UAV detection dataset with multiscale UAVs and complex backgrounds, created using PIC, transforming trajectory data into high-fidelity bounding boxes, enabling automated, precise spatial annotations.
  2. $2$
    Provide the experiments validating PIC’s efficiency in terms of IoU and runtime, along with a comprehensive benchmark of UAVDB using SOTA YOLO-series detectors, including YOLOv8 [11], YOLOv9 [24], YOLOv10 [23], YOLO11 [10], and YOLOv12 [22].
2 Related Work
2.1 Object Detection by Points
Recent studies have explored point-based supervision as a cost-effective alternative to fully annotated datasets for weakly supervised object detection and instance segmentation. These approaches utilize sparse point annotations rather than full bounding boxes or masks, reducing labeling effort while guiding model learning. As shown in [3, 28], a hybrid supervision strategy combines a small subset of fully annotated images with point-labeled images, training a point-to-box regression model to infer bounding boxes. Similarly, [12] introduces a point-guided mask representation, refining object boundaries using minimal point annotations to improve segmentation accuracy while reducing annotation costs. While point-based methods reduce labeling requirements, they face notable limitations. First, they require fine-tuning on domain-specific datasets, making them impractical for dynamic environments with shifting data distributions. Second, training-based optimization incurs considerable computational overhead, restricting their feasibility for large-scale or real-time applications. Third, weak supervision introduces spatial ambiguity, often resulting in imprecise bounding boxes, especially when object boundaries are poorly defined. These challenges underscore the need for a scalable and training-free strategy.
2.2 Bounding Box Extraction via Segmentation
Figure 2: Comparison of bounding box extraction methods across various datasets and cameras. The rightmost column shows our PIC results, which generate high-fidelity bounding box annotations. Other columns depict results from fixed-size bounding boxes, image thresholding [1], GrabCut [20], and SAM [13]. In the last three rows, when the UAV is tiny, or the background is complex, our method remains robust, successfully extracting accurate bounding boxes even in challenging scenarios.
Since learning-based approaches induce some inconvenience, we focus on an out-of-the-box approach to generate the bounding box annotations. As shown in Fig. 1, the goal is to extract high-fidelity bounding boxes for UAVs of varying sizes in videos only with trajectory data. A simple approach assigns a fixed bounding box around the trajectory point, but this lacks flexibility in adjusting box sizes. A more refined alternative segments the fixed size and defines the bounding box using the upper-left and lower-right corners. Image thresholding, as described in [1], is a common technique but becomes ineffective when the contrast between the UAV and background is unclear, requiring manual adjustments. Alternatively, the GrabCut algorithm [20] provides better bounding box accuracy but is computationally expensive and inefficient. Deep learning-based methods, such as DeepGrabCut [27], also demand significant computational resources. Even SOTA models like the Segment Anything Model (SAM) [13] with point prompts encounter domain-specific challenges, resulting in poor segmentation. Fig. 2 illustrates bounding boxes extracted by various methods, with a light gray (#e7e6e6 color hex) background for clearer visualization.
3 Methodology
This section presents the PIC algorithm, details the annotation process, and introduces the UAVDB dataset.
3.1 Patch Intensity Convergence (PIC)
The PIC technique extracts UAV bounding boxes from trajectory annotations via an adaptive inward-outward expansion, ensuring efficient localization without relying on external models or predefined dimensions. The process consists of four steps: initialization, iterative expansion, patch intensity calculation, and convergence assessment.
Figure 3: Stepwise demonstration of the PIC technique applied across various datasets and cameras. The middle columns illustrate the iterative bounding box expansion centered on the UAV, with corresponding pixel intensity values. The rightmost column presents the final PIC annotations along with UAV size and aspect ratio in each scenario.
3.1.1 Initialization
Given a trajectory point $(x_{0},y_{0})$, the bounding box is initialized as a square region $B_{0}$ of size $w_{0}\times h_{0}$:
$B_{0}$ =$=\{(x,y)\mid x_{0}-w_{0}/2\leq x\leq x_{0}+w_{0}/2,$
$\hskip 40.68744pt\quad y_{0}-h_{0}/2\leq y\leq y_{0}+h_{0}/2\}.$
3.1.2 Iterative Expansion
At each step $t$, the bounding box expands outward by a fixed size $\delta$ in all directions:
$w_{t+1}=w_{t}+\delta,\quad h_{t+1}=h_{t}+\delta,\quad t=0,1,\ldots$
The expanded region $B_{t+1}$ captures a progressively larger area around the trajectory point.
3.1.3 Patch Intensity Calculation
The mean pixel intensity at each step inside the bounding box is computed as:
$\mu_{t}=\frac{1}{|B_{t}|}\sum_{(x,y)\in B_{t}}I(x,y).$
where $I(x,y)$ denotes the pixel intensity at $(x,y)$.
3.1.4 Convergence Assessment
Expansion halts when the intensity change between consecutive iterations falls below a threshold $\epsilon$:
$|\mu_{t+1}-\mu_{t}|<\epsilon.$
This criterion ensures that further expansion does not significantly contribute to capturing UAV-relevant pixels, marking the final bounding box boundary.
We apply the PIC technique to the videos and trajectory data from [15], using an initial patch size of $w_{0}=h_{0}=8$ pixels, an expansion step of $\delta=5$ pixels, and a convergence threshold of $\epsilon=4$. For UAVDB, we extract one frame per ten frames (around 10% of the footage) from Tab. 1 to construct the database. The resulting dataset consists of 10,763 training images, 2,720 validation images, and 4,578 test images, as detailed in Tab. 2. Since Dataset 5 in [15] lacks 2D trajectory information, we serve as an unseen scenario, with its detection results presented in the experimental section. Notably, our framework allows flexible adjustment of the extraction rate to generate larger or smaller datasets.
[…]
Table 2: Overview of the UAVDB constructed using the proposed PIC approach. The table shows the distribution of images across different datasets and camera configurations, specifying the number of images used for training, validation, and testing.
Fig. 3 illustrates the stepwise expansion of PIC across different datasets, demonstrating its precision in challenging scenarios. The middle columns depict the incremental bounding box expansions with corresponding pixel intensity values. The rightmost column shows a reference image indicating UAV size as a percentage of the total image area. PIC accurately localizes UAVs across scales, from large (53$\times$52 pixels around 0.133% of the image) to tiny (13$\times$13 pixels around 0.008% of the image), providing the comprehensive and high-fidelity bounding box annotations.
4 Experimental Results
We first evaluate the effectiveness of the proposed PIC approach in terms of IoU metrics and runtime efficiency compared to other methods. We then present comprehensive benchmark results on UAVDB.
4.1 Annotation Accuracy and Runtime Efficiency
Here, human-labeled bounding boxes serve as ground truth annotations. For fixed-size and thresholding [1] approaches, we use a 50$\times$50 region and set the threshold to 150 based on empirical tuning for optimal performance. GrabCut [20] and SAM [13] using the OpenCV package and the ViT-B pre-trained model, respectively. As shown in Tab. 3, the PIC approach achieves the highest IoU while maintaining a minimal runtime of 0.007 seconds, comparable to the fixed-size method. This demonstrates that the computational time of the PIC process is negligible compared to image reading and output. In contrast, human labeling takes an average of 19 seconds per annotation, making it impractical for large datasets with tiny objects. Moreover, despite its advanced segmentation capabilities, SAM struggles with UAV-specific challenges, resulting in the lowest IoU. This illustrates that SAM cannot generalize effectively without retraining on a specific dataset. These results highlight the effectiveness of PIC in providing both accurate and computationally efficient UAV bounding box extraction, making it ideal for large-scale and real-time applications.
Methods Average IoU Runtime human-labeled 1.000 19.00 Fixed-size 0.278 0.007 Thresholding [1] 0.316 0.009 GrabCut [20] 0.425 2.423 SAM [13] 0.249 0.484 PIC (ours) 0.464 0.007
Table 3: Comparison of different UAV bounding box extraction methods regarding average IoU and runtime (seconds).
4.2 Benchmark on UAVDB
We examine YOLOv8 [11], YOLOv9 [24], YOLOv10 [23], YOLO11 [10], and YOLOv12 [22] to benchmark the proposed UAVDB. The experiments were conducted on a high-performance computing (HPC) system [16], utilizing an NVIDIA A100 GPU with 80 GB of memory. All models were trained with an image size of 640, a batch size of 32, 100 epochs, and eight workers. Mosaic augmentation was applied throughout training, excluding the final 10 epochs. Additionally, we fine-tuned the models using officially released pre-trained weights. Tab. 4 summarizes the training time, inference time, model parameters, FLOPs, and average precision (AP) for both validation and test sets. Further, the performance of each model on the validation set across epochs is illustrated in Fig. 4.
Model Training Time Inference Time #Param. (M) FLOPs (G) $\text{AP}^{val}_{50}$ $\text{AP}^{val}_{50-95}$ $\text{AP}^{test}_{50}$ $\text{AP}^{test}_{50-95}$ (hours:mins:sec) (per image, ms) YOLOv8n 01:40:31 0.9 2.685 6.8 0.829 0.522 0.789 0.450 YOLOv8s 01:55:05 1.2 9.828 23.3 0.814 0.545 0.796 0.450 YOLOv8m 02:43:08 1.8 23.203 67.4 0.809 0.538 0.827 0.526 YOLOv8l 03:54:44 2.6 39.434 145.2 0.830 0.563 0.836 0.544 YOLOv8x 04:33:08 3.5 61.597 226.7 0.820 0.554 0.728 0.448 YOLOv9t 02:53:11 2.5 2.617 10.7 0.839 0.501 0.848 0.508 YOLOv9s 03:05:02 2.6 9.598 38.7 0.819 0.517 0.834 0.484 YOLOv9m 05:08:28 4.1 32.553 130.7 0.840 0.507 0.858 0.522 YOLOv9c 06:17:08 5.3 50.698 236.6 0.851 0.544 0.851 0.504 YOLOv9e 08:00:05 6.6 68.548 240.7 0.755 0.414 0.768 0.383 YOLOv10n 02:05:39 0.7 2.695 8.2 0.764 0.492 0.731 0.417 YOLOv10s 02:23:03 1.2 8.036 24.4 0.817 0.530 0.823 0.516 YOLOv10m 03:06:59 1.8 16.452 63.4 0.798 0.531 0.821 0.536 YOLOv10b 03:29:18 2.1 20.413 97.9 0.801 0.517 0.760 0.467 YOLOv10l 04:04:22 2.5 25.718 126.3 0.774 0.502 0.842 0.517 YOLOv10x 05:14:07 3.5 31.586 169.8 0.771 0.507 0.693 0.431 YOLO11n 01:50:00 0.9 2.582 6.3 0.847 0.527 0.856 0.539 YOLO11s 02:07:01 1.2 9.413 21.3 0.826 0.553 0.885 0.578 YOLO11m 03:07:40 1.9 20.031 67.6 0.827 0.588 0.843 0.578 YOLO11l 04:09:45 2.4 25.280 86.6 0.810 0.555 0.798 0.517 YOLO11x 05:20:38 3.6 56.828 194.4 0.812 0.560 0.782 0.534 YOLOv12n 02:15:38 1.8 2.557 6.3 0.857 0.544 0.848 0.531 YOLOv12s 02:44:29 2.0 9.231 21.2 0.869 0.566 0.882 0.565 YOLOv12m 03:34:36 2.6 20.106 67.1 0.866 0.567 0.886 0.584 YOLOv12l 05:10:15 3.1 26.340 88.5 0.870 0.584 0.875 0.590 YOLOv12x 06:35:47 3.9 59.045 198.5 0.879 0.576 0.896 0.569
Table 4: Performance of YOLOv8 [11], YOLOv9 [24], YOLOv10 [23], YOLO11 [10], and YOLOv12 [22] models trained on UAVDB.
Fig. 5 presents YOLO11s, the model achieves the best balance between precision and speed, predictions on Dataset 5, where scenarios were absent from the training data, demonstrating its ability to handle unseen situations. The detection results closely match the UAV sizes, validating the high fidelity of the bounding box annotations in UAVDB. Incorporating these high-quality predicted bounding boxes into the training set can further enhance the model’s capability to detect UAVs.
Figure 4: Performance of YOLOv8 [11], YOLOv9 [24], YOLOv10 [23], YOLO11 [10], and YOLOv12 [22] on validation set.
Figure 5: Detection results predicted by YOLO11s on unseen scenarios.
4.3 Discussion
The proposed PIC generates bounding box annotations with the highest IoU while being approximately 2700$\times$ faster than human labeling. Despite this, the UAVDB remains adequate for training detectors, as shown in Fig. 5. Although the PIC method performs well on current datasets, we recognize that low-altitude UAV flights, with cluttered and rapidly changing backgrounds, may pose challenges. In such dynamic environments, local intensity variations could affect the pixel intensity metric for expanding the bounding box. However, the adaptive nature of PIC, focusing on local intensity changes, allows it to handle moderate variations in background texture. Further improvements for highly dynamic scenarios could include incorporating multi-frame temporal information or background subtraction to enhance robustness and maintain consistent performance.
5 Conclusion
In this paper, we introduced UAVDB, a dataset with precise bounding box annotations facilitated by our proposed PIC technique. PIC offers an intuitive, efficient, and innovative approach to spatial annotation, eliminating the need for manual labeling. UAVDB addresses critical limitations in existing UAV datasets, such as imprecise annotations and limited environmental diversity, significantly improving the applicability of detection algorithms in real-world scenarios. Through IoU and runtime evaluations for PIC and benchmarking with YOLO-series detectors on UAVDB, we demonstrated the versatility of both UAVDB and PIC approaches under varied conditions. These results establish UAVDB as a valuable resource for advancing UAV detection technologies. Looking ahead, PIC’s adaptability opens promising directions for future research. Its lightweight design could be further optimized by incorporating multi-frame temporal information or background subtraction to improve robustness in dynamic environments. Moreover, its flexibility allows fine-tuning for specific domains, ensuring scalability across various UAV detection applications. As UAV detection technology evolves, UAVDB and PIC provide a solid foundation for advancing real-time, large-scale UAV detection in diverse environments.
6 Acknowledgments
We thank the Photogrammetry and Remote Sensing group at ETH Zurich for their invaluable contribution to the multi-view drone tracking dataset [15], which features high-resolution UAV video and 2D trajectory data.